OpenStudy (anonymous):

The rates at which Toni's Tiles and Bob's Bathrooms tile are to one another. The rates at which Toni's Tiles and Rhonda's Restroom Redos tile are to one another. The rates at which Bob's Bathrooms and Rhonda's Restroom Redos tile are to one another. Two items are in a proportional relationship if they the same unit rate. Submit

OpenStudy (anonymous):

Duke wants to hire someone to re-tile his bathroom. The research he found for three local tilers is presented in the table below. He was able to find the average area of their tiling jobs and the time it took the tilers to complete the job. Tiler Area Tiled (square feet) Time (hours:minutes) Toni's Tiles 803 2:12 Bob's Bathrooms 1,460 4:00 Rhonda's Restroom Redos 753 1:30

OpenStudy (anonymous):

HELP MEEEEE

OpenStudy (anonymous):

Calculate the ratio of tiles per minute for each tiler. \[\left\{\frac{803}{132},\frac{1460}{240},\frac{753}{90}\right\} \]Convert the fractions to decimal numbers and then compare.